# Theory and Design of Space-Time Metallic Metasurfaces for Wireless Communications

### v1

@@ Abstract
Index Terms:
Space-time, time-modulation, metasurfaces, analytical method, wireless communications, beamforming, frequency mixing.
@@ I Introduction
A conventional metasurface can be physically described as a thin screen with planar geometry exhibiting spatial modulation [3]. Thanks to this modulation, metasurfaces become excellent terminals for a wide range of applications, such as microwave absorbers [4], beam shaping [5], or lenses [6], among others. Currently, researchers are paying close attention to a new evolution of metasurfaces, the so-called space-time metasurfaces, in which time is employed as an additional modulation to the existing purely spatial one [7, 8]. This introduces additional degrees of freedom as well as richer physics on the system [9, 10, 11]. Likewise, time modulation also enhances the capabilities of the terminals, leading to promising concepts such as intelligent programmable metasurfaces [12] that could be combined and integrated into wireless communications and radars systems [13].
@@
Many of these applications are usually simulated using numerical techniques such as the finite-difference time-domain (FDTD) method [39]. However, the high computational costs and long simulation times that this code entails, have prompted the development of adaptive mesh-based solutions to enhance computational efficiency, such as the Discontinuous Galerkin Time-Domain (DGTD) method [40]. Others numerical solutions based on generalized sheet transition conditions (GSTCs) [41, 42], modal techniques [43] and integral-equation methods [44] show high performance when simulating time-varying metasurfaces. Nonetheless, many times the researchers also search approaches that can provide a deep physical insight of the studied problems. Thus, some alternatives based on analytical and semi-analytical techniques have been reported in the literature [45, 46, 47, 48].
Following this line, fully-analytical circuits for 1D-spatial, 2D-spatial, 3D-spatial [49, 50, 51] and 1D-temporal [52, 53] modulations have been previously reported by the authors. Furthermore, we have recently exploited the use of Lorentz transformations to reduce a (1+1)-D spatiotemporal problem to a time-invariant (1D-spatial) system [54]. Now, in this work, we extend the circuital approach to a different set of (1+1)-D space-time scenarios. The scattering parameters, as well as information on the diffraction spectrum, are extracted in significantly shorter computation times compared to general numerical methods. In fact, the main advantage that this method presents is the possibility to simulate a wideband range of frequencies in a few seconds.
@@ II Analytical Framework
II-A Space-Time Metasurface
Fig. 1 depicts the spatio-temporal system studied in this manuscript. The system is formed by a space-time-modulated diffractive metasurface that manipulates the incident transverse electric (TE) or transverse magnetic (TM) plane waves. The diffractive metasurface is based on different slabs painted in red and blue that alternate their electromagnetic properties in time. Each of the slabs can either become electromagnetically transparent or behave like a solid metal as it is assumed in [55].
Depending on the electromagnetic behavior of each of the slabs, the space-time-modulated diffractive metasurface can operate in three different states: “air” (A), “conductor” (C) and “grating” (G). In the air state, all slabs appears to vanish from an electromagnetic perspective, so full transmission of the incident wave is expected. In the conductor state, all slabs behave as a solid metal achieving a solid metallic wall, so full reflection of the incident wave is expected. In the case of the grating state, the red slabs behave as solid metals, and the blue slabs become electromagnetically transparent, causing the space-time metasurface to operate as a spatially-modulated metallic diffraction grating.
@@ II-A Space-Time Metasurface
II-B Floquet-Bloch Expansion
@@
The previous derivation can easily adapted for TM incidence. In the TM case, the electric and magnetic field orientations are the opposite to the TE case, thus
$\mathbf{E}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{x}}\cdot\mathbf{E}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\cdot\hat{\mathbf{y}}$ (10)
$\mathbf{H}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{y}}\cdot\mathbf{H}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\cdot\hat{\mathbf{x}}\,.$ (11)
@@
II-C Nature of the Space-Time Harmonics
@@
In a simpler scenario where the surrounding media are air ($\varepsilon_{r}=\mu_{r}=1$) and the waves impinge normally ($\theta_{0}=0$) to the space-time metasurface, eq. (II-C) reduces to:
@@
II-D Diffraction Angles and Beamforming
@@
II-E Reflection/Transmission Coefficients
The derivation of the scattering parameters, i.e., the reflection and transmission coefficients, require of prior knowledge of the electric-field profiles at the space-time discontinuity. This is an insightful technique that has been successfully applied in the past for the analysis of 1D-spatial [66, 49], 1D-time [52, 53], 2D-spatial [67, 68, 69] and 3D-spatial [70] periodic systems. Formally, the space-time profile, also referred to as base function (bf), can be described by the function $\mathbf{E}_{\mathrm{bf}}(y,t)$ depending on space and time.
@@ II-E Reflection/Transmission Coefficients
$\mathbf{E}_{t}^{\text{TE/TM},(1)}(y,0,t)=\mathbf{E}_{t}^{\text{TE/TM},(2)}(y,0,t)=\mathbf{E}_{\mathrm{bf}}(y,t)\,,$ (16)
@@
$\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})=\displaystyle\int_{-\frac{P}{2}}^{\frac{P}{2}}\displaystyle\int_{0}^{T_{\text{M}}}E_{\mathrm{bf}}(y,t)\,\text{e}^{-\text{j}(\omega_{n}t+k_{m}y)}\text{d}y\,\text{d}t\,.$ (20)
Note that ${E}_{\mathrm{bf}}(y,t)$ is simply the instantaneous amplitude of the vector $\mathbf{E}_{\mathrm{bf}}(y,t)$ representing the base function and $\widetilde{E}_{\mathrm{bf}}(\omega_{n},k_{m})$ is its Fourier transform.
@@
$(1-R)Y_{00}^{\text{TE/TM},(1)}-(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(1)}\,|N(k_{m},\omega_{n})|^{2}\\ =(1+R)Y_{00}^{\text{TE/TM},(2)}+(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(2)}|N(k_{m},\omega_{n})|^{2}\,,$ (24)
@@ II-F Circuit Topology
Each of the parameters $N(k_{m},\omega_{n})$ taking part in eq. (26) is interpreted as a complex transformer that takes into consideration the coupling between the $(m,n)$-th harmonic and the fundamental ($00$-th) one. Each admittance of each Floquet harmonic is loaded with a complex transformer, except for the fundamental harmonic whose value is $N(k_{0},\omega_{0})=1$. This is sketched in Fig. 3(a). Note that, for the sake of visualization, only a few harmonics of the double infinite sum are represented in Fig. 3(a), and that the transmission lines of the fundamental harmonic are marked in blue.
At a first sight, the circuit topologies shown in Figs. 3(a) and 3(b) coincide with those shown in previous works on space-only and time-only metamaterial configurations [71, 52]. However, the value of the forming admittances and complex transformers, which rule the diffraction behavior of the metasurface, are completely different. In (three-dimensional) space-only systems, the complex transformers are a function of the transverse wavenumbers $k_{m},k_{n}$ and $k_{l}$. In time-only problems, those are a function of the angular frequency $\omega_{n}$. In spatiotemporal systems, the complex transformers are a function of the transverse wavenumbers plus the angular frequency.
@@ II-G Basis Functions
As detailed in Section II-E, the knowledge of the tangential electric field profile $\mathbf{E}_{\mathrm{bf}}(y,t)$ is crucial for the analysis of the space-time metasurface via the analytical method. In this subsection, we describe the analytical expressions for the space-time base functions involving the three states under consideration: air (A), conductor (C) and grating (G).
II-G1 Air state (A)
@@
II-G2 Conductor state (C)
In the conductor state, the space-time metasurface turns into an infinitesimally-thin metallic sheet. If losses are neglected, the metallic sheet can be modeled with PEC condictions, so the tangential electric field must be zero at its surface. This leads to the base function
@@ II-G3 Grating (G)
$E_{\text{bf}}(y,t)=A\sin(\omega_{0}t)\begin{cases}\Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{\frac{1}{2}},\quad&\text{TE polarization}\\ \Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{-\frac{1}{2}},\quad&\text{TM polarization}\,,\\ \end{cases}$ (34)
@@
As previously introduced, a wise combination (A-G, C-G) of the three states in a time-periodic scheme can lead to a rich diffraction phenomenology from which beamformers can benefit. Top panel of Fig. 4(a) sketches the scenario where the space-time metasurface alternates between conductor and grating states (C-G case). As shown, the scheme is periodic in space and time, with periods $P$ and $T_{\mathrm{M}}$, respectively. Moreover, the terms $t_{C}$ and $t_{G}$ indicate the time in which the metasurface is in the conductor and grating states. Note that $T_{\mathrm{M}}=t_{C}+t_{G}$ in this case to retain time periodicity. The basis function that models the C-G scenario is plotted at the bottom of Fig. 4(a) for TE and TM incident polarizations. Null regions indicate that the space-time metasurface is in the conductor state [eq. (33)], while $\cap$- and $\cup$-shaped regions [eq. (34)] are associated to the grating state under TE and TM incidences, respectively.
@@
Transients between air, conductor and grating states are neglected along the document. In general, the electric field at the space-time metasurface takes a nonzero time in reaching the steady state for the considered metasurface state. However, transient phenomena has no appreciable impact on the diffraction pattern as long as the considered operating frequencies are sufficiently low. In our particular case, we must ensure operation far below the plasma frequency $\omega_{p}$ of the metal under consideration and time periods ($T_{0}$ and $T_{\text{M}}$) much greater than the relaxation time $\tau$ of the free charges. Metals that are good conductors such as copper, aluminum, silver or gold have plasma frequencies of the order of $\omega_{p}\sim 10^{15}$ rad/s and relaxation times $\tau\sim 10^{-14}$ s. Thus, transient times are several orders of magnitude shorter than the periods of the incident wave and time modulation in the radio, microwave and millimeter-wave frequencies. In practice, the operation frequency of the analytical approach is limited down to the low-THz regime.
@@
The considered basis functions also limit the frequency range of the analytical approach. Specifically, the most constraining basis function is the one that models the grating state. It was reported in previous works on spatially-periodic metasurfaces that the maximum operating frequency is limited, from a practical perspective, to $f_{\text{max}}\sim 1.5c/P$ [74, 75]. Thus, the proposed analytical works in a wide frequency range, even within the grating lobe regime. In the case of considering oblique incidence, the analytical approach is expected to work fine as well for both TM and TE polarization states, but the maximum operating frequency could reduce to $f_{\text{max}}\sim c/P$ for large incident angles. Moreover, the slit width should be less than $W\lesssim 0.7P$ to obtain accurate results. That is, the analytical approach is wideband, robust to large incidence angles and not strictly constrained to the subwavelength regime. Taking into account the aforementioned limitations, from an engineering perspective, we will be able to excite many propagative harmonics if we go into the diffraction regime with $P\gg w$. This scenario would lead to a richer diffraction pattern.

### v2

# Space-time Metallic Metasurfaces for Frequency Conversion and Beamforming

@@ I Introduction
A conventional metasurface can be physically described as a thin screen with planar geometry exhibiting spatial modulation [3]. Thanks to this modulation, metasurfaces become excellent terminals for a wide range of applications, such as microwave absorbers [4], beam shaping [5], or lenses [6], among others. Currently, researchers are paying close attention to a new evolution of metasurfaces, the so-called space-time metasurfaces, in which time is employed as an additional modulation to the existing purely spatial one [7, 8]. This introduces additional degrees of freedom as well as richer physics on the system [9, 10, 11]. Likewise, time modulation also enhances the capabilities of the active terminals [12], leading to promising concepts such as intelligent programmable metasurfaces [13] that could be combined and integrated into wireless communications and radars systems [14].
@@
Unfortunately, the lack of available commercial solutions to simulate complex behavior in space-time-modulated media hampers the analysis and design of these metadevices. At present, commercial simulator COMSOL Multiphysics is the only direct alternative for simulating spatiotemporal metamaterials. Nonetheless, it naturally presents the limitations inherent to a general-purpose electromagnetic simulator. As an in-house alternative, many of these applications are usually simulated using numerical techniques such as the finite-difference time-domain (FDTD) method [39]. However, the high computational costs and long simulation times that this code entails, have prompted the development of adaptive mesh-based solutions to enhance computational efficiency, such as the Discontinuous Galerkin Time-Domain (DGTD) method [40]. Others numerical solutions based on generalized sheet transition conditions (GSTCs) [41, 42], modal techniques [43] and integral-equation methods [44] show high performance when simulating time-varying metasurfaces. Nonetheless, many times the researchers also search approaches that can provide a deep physical insight of the studied problems. Thus, some alternatives based on analytical and semi-analytical techniques have already been reported [45, 46, 47, 48].
In this work, we implement an analytical technique for the analysis of space-time metallic metasurfaces that can commute between three states: “air”, “conductor” and “grating”. As it will be show, a wise temporal combination of the three mentioned spatial states can produce a rich diffraction spectrum. Transmission and reflection levels can be tuned at convenience depending on the amount of time the metasurface is in these states. The air state favours transmission, while the conductor state favours reflection. The addition of the grating state adds richness and variety to the diffraction spectrum. The present space-time metasurface is of potential interest for related applications such as frequency conversion and beamforming.
Our analytical proposal is based on a Floquet-Bloch series expansion of the electromagnetic fields and integral-equation techniques that account for the space-time boundary conditions. The analytical formulation comes with an equivalent circuit that gives phyisical insight on spatiotemporal diffraction. Interestingly, the associated circuit model comes from first-principle computations; namely, the topology and elements of the circuit are not imposed by the designer, but arise as a consequence of the analytical formulation itself. Analytical circuits for spatially-modulated [49, 50, 51] and temporally-modulated [52, 53, 54] systems have been previously reported by the authors. The analytical circuit is a physically-insightful object that will give us information on the nature of the space-time harmonics, coupling between them, preferred diffraction directions and scattering parameters for both transverse magnetic (TM) and transverse electric (TE) polarizations. Moreover, computational times are minimal: a wideband range of frequencies can be simulated in less than a second.
@@ II Analytical Framework
II.1 Space-Time Metasurface
Fig. 1 depicts the spatio-temporal system studied in this manuscript. The system is formed by a space-time-modulated diffractive metasurface that manipulates the incident transverse electric (TE) or transverse magnetic (TM) plane waves. The diffractive metasurface is based on different strips painted in red and blue that alternate their electromagnetic properties in time. Each strip can either become electromagnetically transparent or behave like a solid metal as it is assumed in [55].
Depending on the electromagnetic behavior of each of the strips, the space-time-modulated diffractive metasurface can operate in three different states: “air” (A), “conductor” (C) and “grating” (G). In the air state, all strips appear to vanish from an electromagnetic perspective, so full transmission of the incident wave is expected. In the conductor state, all strips behave as a solid metal achieving a solid metallic wall, so full reflection of the incident wave is expected. In the case of the grating state, the red strips behave as solid metals, and the blue strips become electromagnetically transparent, causing the space-time metasurface to operate as a spatially-modulated metallic diffraction grating.
@@ II.1 Space-Time Metasurface
II.2 Floquet-Bloch Expansion
@@
The previous derivation can be easily adapted for TM incidence. In the TM case, the electric and magnetic field orientations are the opposite to the TE case, thus
$\mathbf{E}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{x}}\cdot\mathbf{E}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\hat{\mathbf{y}}$ (10)
$\mathbf{H}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{y}}\cdot\mathbf{H}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\hat{\mathbf{x}}\,.$ (11)
@@
II.3 Nature of the Space-Time Harmonics
@@
In a simpler scenario where the surrounding media are air ($\varepsilon_{r}=\mu_{r}=1$) and the waves impinge normally ($\theta_{0}=0$) to the space-time metasurface, eq. (13) reduces to:
@@
II.4 Diffraction Angles and Beamforming
@@
II.5 Reflection/Transmission Coefficients
The derivation of the scattering parameters, i.e., the reflection and transmission coefficients, require of prior knowledge of the electric-field profiles at the space-time discontinuity. Formally, the space-time profile, also referred to as basis function (bf), can be described by the function $\mathbf{E}_{\mathrm{bf}}(y,t)$ depending on space and time.
@@ II.5 Reflection/Transmission Coefficients
$\mathbf{E}_{t}^{\text{TE/TM},(1)}(y,0,t)=\mathbf{E}_{\mathrm{bf}}(y,t)=\mathbf{E}_{t}^{\text{TE/TM},(2)}(y,0,t)\,,$ (16)
@@
$\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})=\displaystyle\int_{-\frac{P}{2}}^{\frac{P}{2}}\displaystyle\int_{0}^{T_{\text{M}}}E_{\mathrm{bf}}(y,t)\,\text{e}^{-\text{j}(\omega_{n}t-k_{m}y)}\text{d}y\,\text{d}t\,.$ (20)
The term ${E}_{\mathrm{bf}}(y,t)$ is simply the instantaneous amplitude of the vector $\mathbf{E}_{\mathrm{bf}}(y,t)$ representing the base function and $\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})$ is its Fourier transform.
The imposition of the boundary conditions in eq. (16) presents a fundamental difference with respect to other analytical methods that are also based on the use of Bloch-Floquet series [42]. Other methods, after expanding the fields in Bloch-Floquet series to the left (1) and right (2) sides of the space-time interface, directly impose continuity between them, i.e., $\mathbf{E}_{t}^{\text{TE/TM},(1)}=\mathbf{E}_{t}^{\text{TE/TM},(2)}$. Then, the fields are projected on each Floquet harmonic to reach a linear system of equations with complex values that gives the Floquet coefficients $E_{mn}$. This technique share similarities with mode-matching methods. In our case, the approach is different. We project separately the fields of the left and right sides of the space-time discontinuity onto a single basis function $E_{\mathrm{bf}}$ of our choice, i.e., $\mathbf{E}_{t}^{\text{TE/TM},(1)}=\mathbf{E}_{\mathrm{bf}}$ and $\mathbf{E}_{t}^{\text{TE/TM},(2)}=\mathbf{E}_{\mathrm{bf}}$. In this case, instead of reaching a complex-valued linear system of equations, we directly obtain the value of the Floquet coefficients $E_{mn}$. This approximation has been successfully applied in the past for the analysis of 1D-spatial [49, 65], 1D-time [52, 53], 2D-spatial [66, 67, 68] and 3D-spatial [51] periodic systems. It significantly reduces the computation time by avoiding solving the system of equations and allows to extract the associated equivalent circuit shown in Fig. 3(a).
We project separately the fields of the left and right sides of the space-time discontinuity onto a single basis function $E_{\mathrm{bf}}$ of our choice, i.e., $E^{(1)}=E_{\mathrm{bf}}$ and $E^{(2)}=E_{\mathrm{bf}}$ [see eq. (16)].
@@
$(1-R)Y_{00}^{\text{TE/TM},(1)}-(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(1)}\,|N(k_{m},\omega_{n})|^{2}\\ =(1+R)Y_{00}^{\text{TE/TM},(2)}+(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(2)}|N(k_{m},\omega_{n})|^{2}\,,\$ (24)
@@ II.6 Circuit Topology
Each of the parameters $N(k_{m},\omega_{n})$ taking part in eq. (26) is interpreted as a complex transformer that takes into consideration the coupling between the $(m,n)$-th harmonic and the fundamental ($0,0$)th one. Each admittance of each Floquet harmonic is loaded with a complex transformer, except for the fundamental harmonic whose value is $N(k_{0},\omega_{0})=1$. This is sketched in Fig. 3(a). Note that, for the sake of visualization, only a few harmonics of the double infinite sum are represented in Fig. 3(a), and that the transmission lines of the fundamental harmonic are marked in blue.
At a first sight, the circuit topologies shown in Figs. 3(a) and 3(b) coincide with those shown in previous works on space-only [69] and time-only [52] metamaterial configurations. However, the value of the forming admittances and complex transformers, which rule the diffraction behavior of the metasurface, are completely different. In (three-dimensional) space-only systems, the complex transformers are a function of the transverse wavenumbers $k_{m},k_{n}$ and $k_{l}$. In time-only problems, those are a function of the angular frequency $\omega_{n}$. In spatiotemporal systems, the complex transformers are a function of the transverse wavenumbers plus the angular frequency.
@@ II.7 Basis Functions
As detailed in Section II.5, the knowledge of the tangential electric field profile (basis function) $\mathbf{E}_{\mathrm{bf}}(y,t)$ is crucial for the analysis of the space-time metasurface via the analytical method. We expect to obtain accurate results as long as the selected basis functions carry actual physical information of the electromagnetic fields taking place in the considered scenario. In this subsection, we describe the analytical expressions for the space-time base functions involving the three states under consideration: air (A), conductor (C) and grating (G).
II.7.1 Air state (A)
@@
II.7.2 Conductor state (C)
In the conductor state, the space-time metasurface turns into an infinitesimally-thin metallic sheet. If losses are neglected, the metallic sheet can be modeled with PEC conditions, so the tangential electric field must be zero at its surface. This leads to the base function
@@ II.7.3 Grating (G)
$E_{\text{bf}}(y,t)=A\sin(\omega_{0}t)\begin{cases}\Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{\frac{1}{2}},\quad&\text{TE pol.}\\ \Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{-\frac{1}{2}},\quad&\text{TM pol.}\,,\\ \end{cases}$ (34)
@@
As previously introduced, a wise combination (A-G, C-G) of the three states in a time-periodic scheme can lead to a rich diffraction phenomenology from which beamformers can benefit. Left side of Fig. 4(a) sketches the scenario where the space-time metasurface alternates between conductor and grating states (C-G case). As shown, the scheme is periodic in space and time, with periods $P$ and $T_{\mathrm{M}}$, respectively. Moreover, the terms $t_{C}$ and $t_{G}$ indicate the time in which the metasurface is in the conductor and grating states. Note that $T_{\mathrm{M}}=t_{C}+t_{G}$ in this case to retain time periodicity. The basis function that models the C-G scenario is plotted at the rigth side of Fig. 4(a) for TE and TM incident polarizations. Null regions indicate that the space-time metasurface is in the conductor state [eq. (33)], while $\cap$- and $\cup$-shaped regions [eq. (34)] are associated to the grating state under TE and TM incidences, respectively.
@@
II.8 On the Limitations of the Analytical Approach
As we fundamentally work with air and metal states, we first have to define the properties of these elements. Without loss of generality, the air state is modeled as a dielectric of electrical properties $\varepsilon_{r}=\mu_{r}=1$. A first approximation comes when defining the properties of metal-based states: conductor and grating. We consider here that metals are modeled as perfect electric conductors (PEC). The assumption of PEC is expected to give accurate results in radio, microwave and millimeter-wave frequencies, where good conducting metals such as copper, gold or silver behave essentially as PEC. When working with metasurfaces and frequency selective surfaces (FSS), it is a common practice in the literature to approximate the response of metals as perfect electric conductors in the aforementioned frequency ranges [49, 66, 71]. In fact, we expect that the PEC assumption give accurate results up to the low-THz regime ($<10$ THz), as long as transients can be neglected in the system. However, the approximation of PEC to model metal-based states would cease to be physical, and therefore valid, in the optical range and at higher frequencies.
The considered basis functions also limit the frequency range of the analytical approach. Of the three states in which the space-time metasurface can be (air, conductor and grating), the grating state is the most limiting from an analytical perspective, but the richest one in terms of diffraction. In [71], it is reported that the maximum operating frequency to be covered with eq. (34) may go beyond to $f\sim 1.5c/P$. For instance, for two different spatial periods of $P=3$ mm and $P=3$ $\mu$m, the maximum frequencies in the grating state may not overpass 150 GHz and 150 THz, respectively. When oblique incidence is considered, this range could decrease to $f_{\mathrm{max}}\sim c/P$ (100 GHz and 100 THz in the former example). Moreover, the slit width should be less than $W\lesssim 0.7P$ to obtain accurate results [71]. Otherwise, it would be more convenient to directly work with currents instead of with electric field profiles [49]. As seen, the frequency range is wide, but naturally limited, when modeling the grating state.
The maximum operation frequency $f_{\mathrm{max}}$ of the analytical approach would be determined by the worst-case scenario of the two main approximations discussed above. If the maximum frequency is exceeded, the accuracy in the results is expected to degrade significantly. Above $f_{\mathrm{max}}$, the proposed basis functions could not be representative of the actual physical phenomena occurring in the spatiotemporal metastructure. Additionally, metals could stop being fully-reflective objects (PEC scenario) and start being semi-transparent from the incident wave’s perspective. Therefore, the actual reflection and transmission coefficients can drastically change with respect to the theoretical estimations.
Moreover, above the maximum operation frequency $f_{\mathrm{max}}$ of the analytical circuit, frequency dispersion and transients could play an important role in the response of the metastructure. Transient effects, out of the scope of the present manuscript, can completely redefine the actual electromagnetic response of the space-time metasurface. Transients are expected to be non-negligible at frequencies close to the plasma frequency $\omega_{p}$ (Drude model). In good conductors such as copper, gold or silver, the plasma frequency is of the order of $\omega_{p}\sim 10^{16}$ rad/s, and the relaxation times are $\tau\sim 10^{-14}$ s. Thus, the effect of transients should be appreciable around frequencies of the order of $10^{14}$ rad/s. Below this limit (radio, microwaves, millimeter-waves), transients may be neglected without significantly affecting the response of the system.